Can someone help me with these two problem. I have to find b²-4ac and number of real solutions or graph the problem and I can figure them out. Thank you for your time.
9m²+16=24m
-3x²+7x=0

Answer by mananth(16946)

9m²+16=24m
add -24m to both sides
9m^2-24m+16=24m-24m
9m^2-24m+16=0
x1= (-(-24)+sqrt((-24)^2-4*9*16))) /(2*9)
x1= 1.333
b^2-4ac =(-24)^2-4*9*16)=0
There is only 1 root
-3x²+7x=0
-x(3x-7)=0
x=0
or
3x-7=0
x=7/3